A chemist wants to make a mixture which is 62% iodine. How many liters of 60% iodine should be combined with 9 liters of 66% iodine to create a 62% solution?
Answer by jorel1380(3719)

n(.60)+9(.66)=(n+9).62
.60n+5.94=.62n+5.58
.02n=.36
n=.36/.02=18
You need 18 liters of 60% iodine solution for your mix..